Foam Products, Inc., makes foam seat cushions for the automotive and aerospace industries. The company’s activity-based costing system has four activity cost pools, which are listed below along with their activity measures and activity rates: Activity Cost Pool Activity Measure Activity Rate Supporting direct labor Number of direct labor-hours $ 11 per direct labor-hour Batch processing Number of batches $ 97 per batch Order processing Number of orders $ 288 per order Customer service Number of customers $ 2,612 per customer The company just completed a single order from Interstate Trucking for 2,100 custom seat cushions. The order was produced in two batches. Each seat cushion required 0.7 direct labor-hours. The selling price was $141.30 per unit, the direct materials cost was $102 per unit, and the direct labor cost was $13.70 per unit. This was Interstate Trucking’s only order during the year. Required: Calculate the customer margin on sales to Interstate Trucking for the year.
Answer #1
Interstate Trucking
Customer Margin—ABC Analysis
Sales 296730 =2100*141.30
Costs:
Direct materials 214200 =2100*102
Direct labor 28770 =2100*13.70
Supporting direct labor 16170 =2100*0.7*11
Batch processing 194 =97*2
Order processing 288
Customer service overhead 2612
262234
Customer margin 34496